OpenStudy (anonymous):

Evaluate 8 Σ 3n/2 n=1

OpenStudy (anonymous):

its is 3(8)(9)/4=54

OpenStudy (watchmath):

go sleep brackett :D

OpenStudy (anonymous):

Thank you, brakett!

OpenStudy (anonymous):

:D i wanto get 44 tonight...just 70 short :D

OpenStudy (anonymous):

\[\sum_{k=1}^n k =\frac{n(n+1)}{2}\] so for this one factor out the \[\frac{3}{2}\] to get \[\frac{3}{2}\sum_{k=1}^8k\]

OpenStudy (anonymous):

what is 44?

OpenStudy (anonymous):

sat? in wat context are u asking ?